OpenStudy (anonymous):

Evaluate and write the answer in scientific notation: (4x10^5)/(2.5x10^8) A. 1.6 x 10^5/8 B. 1.6 x 10^-3 C. 1.6 x 10^3

OpenStudy (unklerhaukus):

\[\frac{4\times10^5}{2.5\times10^8}=\frac{4}{2.5} \times 10^{-3}=B.\]

OpenStudy (anonymous):

thankyou!!!